What is the next sequence of numbers; given: 4,18,200,20402
Answer by vleith(2517)

given one term, to find the next one - follow these steps
1) Divide the first number by 2
2) Add one
3) Square
4) Multiply 2
So, given 4:
divide by 2 to get 2
Add 1 to get 3
Square to get 9
Multiply by 2 to get 18
Given 200:
1) 200/2 = 100
2) 100+1 = 101
3) 101^1 = 10201
4) Mult by 2 to get 20402
I know, clunky, but it holds for those 4 numbers anyway